OpenStudy (anonymous):

The sum of three consecutive numbers is 138. What are the three numbers?

OpenStudy (anonymous):

since they are consecutive, any three no's will be a,a+1,a+2; .:3a=138-3=135; a=135/3=45 the no's are 45,46,47.

OpenStudy (anonymous):

Thank you..